# Gradient elasticity in Swift-Hohenberg and phase-field crystal models

### v1

@@ 2.3 Special second-strain gradient elasticity
with $\tau_{ijmn}=\partial f/\partial(\partial_{n}\partial_{m}\varepsilon_{ij})$ the triple stress tensor and $\nabla^{4}$ the bi-Laplacian. Constants relate as $\omega^{2}=\varrho_{1}^{2}+\varrho_{2}^{2}$ and $\gamma^{4}=\varrho_{1}^{2}\varrho_{2}^{2}$. Note that to the leading order ($\gamma=0$), this theory reduces to a single characteristic length $\omega$, i.e. to GE-1 with $\ell_{1}=\ell_{2}=\omega$. This formulation has been exploited to obtain analytical solutions, e.g., for dislocation-induced deformations, which will be used in the following.
3 Swift-Hohenberg energy functional and elasticity
@@ 3.1 Free Energy for the microscopic density field
The Swift-Hohenberg (SH) and Phase-Field Crystal (PFC) model describe the non-conservative and conservative dynamics of a periodic order parameter $\psi\equiv\psi(\mathbf{r})$. The dynamical equation can be written in terms of the variation of a free energy functional $F$ that, in its simplest form, can be written [27, 32]
$F=\int_{\Omega}{\rm d}\mathbf{x}\bigg{[}\frac{1}{2}\psi\mathcal{L}\psi+\frac{1}{4}\psi^{4}\bigg{]},$ (24)
with $\mathcal{L}=(q_{0}^{2}+\nabla^{2})^{2}-\epsilon$. This differential operator realizes an approximation of the first peak in the structure factor (at $q_{0}$) of a periodic structure, namely a stripe or a crystalline phase. The so-called SH equation corresponds to the non-conservative $L^{2}$-gradient flow
@@
Eq. (24) realizes the minimal free energy to model order-disorder phase transition as its minimizing fields are either constant or periodic. As such, its form emerges in different contexts. $F_{\rm\psi}$ can be obtained as the non-dimensional form of the free energy describing liquid-solid transition with the order parameter corresponding to the microscopic density field $\rho$ [33]
$F_{\rm LS}=\int_{\Omega}{\rm d}\mathbf{r}\bigg{[}\frac{1}{2}\rho\bigg{(}U\Delta T+\lambda(q_{0}^{2}+\nabla^{2})^{2}\bigg{)}\rho+\frac{V}{4}\rho^{4}\bigg{]},$ (27)
with
$\mathbf{r}=\mathbf{x}q_{0},\quad\psi=\rho\sqrt{\frac{V}{\lambda q_{0}^{4}}},\quad\epsilon=-\frac{U\Delta T}{\lambda q_{0}^{4}},\quad F_{\psi}=\frac{F_{\rm LS}}{F_{0}},\quad F_{0}=\frac{1}{V}\lambda^{2}q_{0}^{8-d},$ (28)
$\Delta T=(T-T_{\rm s})/T_{\rm s}$, the relative temperature w.r.t the spinodal temperature $T_{\rm s}$, and $U$, $V$, $\lambda$ parameters. An SH energy functional for pure materials may also be derived—upon approximations—from the classical density functional theory [51, 52, 53] as discussed, e.g., in Refs. [54, 55, 34, 56]. One may consider the dimensionless deviation of the density $\rho$ from its average $\bar{\rho}$, $n=(\rho-\bar{\rho})/\bar{\rho}$, as order parameter. The local part, namely the nonlinear polynomial $P(n)$ is then obtained as an approximation of the ideal-gas free energy, expanded as a power series up to the fourth order. The differential operator is obtained via the approximation of the $n$-point correlation $C_{n}(\mathbf{r}_{1},\mathbf{r}_{2},...)$ in the excess energy term. The correlation function is truncated at $C_{2}$ and approximated via an expansion up to the fourth order as $C_{2}(|\mathbf{r}-\mathbf{r}^{\prime}|)=(c_{0}-c_{2}\nabla^{2}+c_{4}\nabla^{4})\delta(\mathbf{r}-\mathbf{r}^{\prime})$. The energy functional thus reads
$\begin{split}F_{c}(\mathbf{r})=&\int_{\Omega}{\rm d}\mathbf{r}\left[P(n(\mathbf{r}))-\frac{1}{2}n(\mathbf{r})\int_{\Omega}C_{2}(|\mathbf{r}-\mathbf{r}^{\prime}|)n(\mathbf{r}^{\prime}){\rm d}\mathbf{r}^{\prime}\right]\\ \approx&\int_{\Omega}{\rm d}\mathbf{r}\left[P(n(\mathbf{r}))-\frac{1}{2}n(\mathbf{r})(c_{0}-c_{2}\nabla^{2}+c_{4}\nabla^{4})n(\mathbf{r})\right],\end{split}$ (29)
and resembles Eq. (24) with, notably, $R=\sqrt{2|c_{4}|/c_{2}}=1/q_{0}$. In the structural PFC model (XPFC) [57, 58], a more detailed control of the crystal structure is obtained by defining directly the correlation function in (29). Such a formulation is particularly well-suited for Fourier pseudo-spectral methods and $C_{2}$ is typically defined via its Fourier transform $\widehat{C}_{2}(\mathbf{k})$.
A multimode SH energy functional has also been proposed [59] with a differential operator approximating $M$ peaks in the structure factor,
@@
$F_{\psi}=\int_{\Omega}{\rm d}\mathbf{x}\bigg{[}A\psi\mathcal{L}_{M}\psi+B\psi^{2}+C\psi^{3}+D\psi^{4}\bigg{]}.$ (31)
@@ 3.2 Amplitude expansion
In the amplitude expansion of the PFC model (APFC) [47], a free energy $F_{\eta}$ dependent on amplitudes can be derived from $F_{\psi}$. For a one-mode approximation of $\psi$ ($M=1$), this has been obtained via a renormalization group approach or by substituting (32) into the expression of the free energy functional and integrating over the unit cell [62, 63, 64]. Here, with the shortest reciprocal-space vector set to $|\mathbf{k}_{n}|=q_{0}=1$ and focusing on bulk systems where $\psi_{0}(\mathbf{r})$ can be assumed constant ($\psi_{0}(\mathbf{r})=\bar{\psi}$), we consider the amplitude approximation of the multimode SH functional (27) introduced in [65]
@@
with $\mathcal{G}_{n}=\nabla^{2}+2{\rm i}\mathbf{k}_{n}\cdot\nabla$, $g^{\rm s}=\frac{B^{\prime}}{2}\zeta_{2}+\frac{C^{\prime}}{3}\zeta_{3}+\frac{D^{\prime}}{4}\zeta_{4}+E^{\prime}$, $\zeta_{p}$ complex polynomials of order $p$ in the amplitudes depending on the lattice symmetry [47], and
$\Gamma_{n}=\frac{\prod_{{q}_{m}\neq|\mathbf{k}_{n}|}[(q_{m}^{2}-|\mathbf{k}_{n}|^{2})^{2}+b_{m}]}{\prod_{{q}_{m}\neq q_{0}}[(q_{m}^{2}-q_{0}^{2})^{2}+b_{m}]}$ (34)
@@
3.3 Elastic energy with strain gradient terms
As the displacement field varies over large length scales, elasticity in the SH or PFC model can be described by looking at amplitudes ($\{\eta_{j}\}$) [40, 47]. In this context, the elastic energy density, namely the part of the energy depending on the displacement $\mathbf{u}$ is
@@ 3.3 Elastic energy with strain gradient terms
Under the simplifying assumption of constant amplitude moduli ($|\nabla\varphi_{n}|=0$),
@@
with $\mathbf{U}$ a non-linear strain tensor and $\mathbf{G}$ encoding gradient elasticity contributions. Then, neglecting higher-order nonlinear terms, we have
$w=A^{\prime}\sum_{n}^{N}\Gamma_{n}\varphi_{n}^{2}\left\{4[(\mathbf{k}_{n})_{i}(\mathbf{k}_{n})_{j}(\partial_{i}u_{j})]^{2}+[(\mathbf{k}_{n})_{i}\nabla^{2}u_{i}]^{2}\right\},$ (38)
with the first and second terms in the sum corresponding to linear and strain-gradient elasticity terms, respectively. The latter may be rewritten as
@@
given that, for each mode (i.e., for each family of equal-length $\mathbf{k}$), it holds true that
$\begin{split}\sum_{n}(\mathbf{k}_{n})_{i}^{2}=Q,\qquad\text{and}\qquad\sum_{n}(\mathbf{k}_{n})_{i}(\mathbf{k}_{n})_{j}=0,\quad\forall i\neq j,\end{split}$ (40)
where $K$ and $Q$ depend on the specific lattice symmetry and the number of modes considered.
To determine gradient-elasticity constants and characteristic lengths we then need to express the quantity (39) in terms of strain components. We consider 2D systems for brevity and without loss of generality. We have that
@@
where we exploited the identity $\partial_{ii}^{2}u_{j}=\partial_{i}(2\varepsilon_{ij}-\partial_{j}u_{i})=2\partial_{i}\varepsilon_{ij}-\partial_{j}\varepsilon_{ii}$. This functional form for the elastic energy density matches the strain-gradient formulation for isotropic media. We may then compare terms of $w$ to the general form of the elastic energy for isotropic materials (12):
@@
For three-dimensional systems, the expressions of coefficient $a_{i}$ are analogous. This follows by considering the additional term $(\nabla^{2}u_{z})^{2}$ on the left-hand side of Eq. (41) and proper extension of the right-hand side. The material length scales, as can be directly deduced from the first and fifth equations in (42), are
@@
with $\ell_{3}=0$. As a result, under the assumptions considered in this section, we establish that the SH energy functional encodes isotropic GE-1. We note that while no centrosymmetric crystal can verify the conditions for isotropic strain-gradient elasticity (due to the Hermann theorem [66]), for cubic materials that are nearly isotropic with respect to the constitutive four-rank tensor $\mathbb{C}$, a Voigt-type averaging procedure may be employed for the sixth-rank constitutive tensor $\mathbb{D}$ in order to compute effective isotropic gradient-elastic constants [67]. In the following, we look at the expressions of the elastic energy density obtained for specific lattice symmetries.
4 Elastic constants and characteristic lengths from the SH functional
4.1 Stripe Phase
Stripe arrangements, e.g., in smectic phases [68, 69, 70, 71] are a class of patterns described by SH and PFC models [27, 32]. The corresponding minimizer of the SH energy functional is well represented by waves with a single reciprocal lattice vector $\mathbf{k}$, namely with $\psi$ described as in Eq. (32) with a single term in the sum. Assuming an arbitrarily oriented wave vector $\mathbf{k}=k_{0}(\cos(\phi),\sin(\phi))$, Eq. (38) reduces to:
$\mathbb{C}=8A^{\prime}\varphi^{2}k_{0}^{4}\begin{pmatrix}\cos^{4}(\phi)&\cos^{2}(\phi)\sin^{2}(\phi)&\cos^{3}(\phi)\sin(\phi)\\ *&\sin^{4}(\phi)&\cos(\phi)\sin^{3}(\phi)\\ *&*&\cos^{2}(\phi)\sin^{2}(\phi)\end{pmatrix}.$ (46)
Notice that in this case, the condition (39) is not met, and strain-gradient constants cannot be simply recast in the characteristic lengths $\ell_{i}$.
4.2 Crystalline phases
The analysis of the density $\psi$ for crystalline phases in SH and PFC models, e.g., to determine phase diagrams and elastic constants, is typically carried out by looking at the smallest number of modes in Eq. (32) required to describe a given lattice symmetry. In the coarse-grained APFC model, the number of modes enter the definition of the variables ($\eta_{n}$) and thus leads to different solutions, i.e., different reconstructed $\psi$ via Eq. (32). We consider here different crystalline phases: triangular, square, bcc, and fcc. For some representative cases, we analyze different numbers of modes. We report the elastic energy density for the different symmetries/approximations, set by specific choices for $\mathbf{k}_{n}$ with $\min_{n=1}^{N}|\mathbf{k}_{n}|=1$ (without loss of generality considering adimensional spatial lengths $\mathbf{x}=\mathbf{r}R^{-1}=\mathbf{r}q_{0}$), the corresponding elastic constants, and GE characteristic lengths for small deformations and neglecting nonlinearities (see Table 1 and Fig. 2). We introduce the scaling factor $k_{0}$ such that $a_{0}=2\pi/k_{0}$ is the interatomic distance.
@@ 4.2 Crystalline phases
The elastic constants and GE characteristic lengths obtained via Eq. (21) from the elastic energy densities above are reported in Table 1. Note that for all lattice symmetries, the constraint $C_{12}=C_{44}$ holds. Interestingly, this corresponds to the Cauchy relation derived in the lattice-theoretical description of constitutive tensors for central force interaction between atoms [72, 67].
@@
We find that $\ell_{2}/\ell_{1}=\sqrt{3}$, as a result of the general form of the (isotropic) strain-gradient terms reported in Sect. 3.3. Typical values found or adopted in the literature for this ratio are close to one [73, 74]. For a one-mode approximation of $\psi$ (one length of $\mathbf{k}$ considered), length scales are constants. For approximations of $\psi$ considering more than one mode, characteristic lengths depend on $\varphi_{i}$. Figure 1 illustrates the bulk free energy density for the tri-2 (Fig. 1a) and fcc-2 (Fig. 1b) case as a function of $\varphi_{1}$ and $\varphi_{2}$ for selected parameters. The global minima of these free energy densities (blue squares) correspond to the equilibrium phase. Values of $\varphi_{1}$ and $\varphi_{2}$ vary with the free energy parameters as dictated by the energy landscape so they are not free parameters entering $\alpha$. An illustration of their variation is reported in Fig. 1c (tri-2) and 1d (fcc-2).
Figure 2 reports the values of $\ell_{1,2}$ by varying $B$ in the free energy for a selected parameter range where all the symmetries can be explored for $B<B_{c}$ and $B_{c}$ corresponding to the order-disorder critical point [47]. Here the effect of changes in $\varphi_{i}$ is then quantified for a specific choice of parameters. We note that a limited change is observed among all the explored ranges ($\lesssim 10\%$). The lattice symmetry is found to affect the values of the GE length scales. At small enough $B$, the number of modes considered does not impact the characteristic lengths whereas a clear deviation is observed at larger $B$ (Fig. 2). In the context of APFC, these approximations realize different models as the choice of modes dictates the variables to solve for in the system (the amplitudes $\eta_{j}$). Instead, for PFC and SH, considering a larger number of modes results in a better estimation of the model’s elastic properties.

### v2

@@ 2.3 Special second strain-gradient elasticity
with $\tau_{ijmn}=\partial f/\partial(\partial_{n}\partial_{m}\varepsilon_{ij})$ the triple stress tensor and $\nabla^{4}$ the bi-Laplacian. Constants relate as $\omega^{2}=\varrho_{1}^{2}+\varrho_{2}^{2}$ and $\gamma^{4}=\varrho_{1}^{2}\varrho_{2}^{2}$. Note that to leading order ($\gamma=0$), this theory reduces to a single characteristic length $\omega$, i.e. to GE-1 with $\ell_{1}=\ell_{2}=\omega$. This formulation has been exploited to obtain analytical solutions, e.g., for dislocation-induced deformations, which will be used in the following.
3 Swift-Hohenberg and phase-field crystal models
@@ 3.1 Free Energy for the microscopic density field
The Swift-Hohenberg (SH) and phase-field crystal (PFC) models describe the nonconservative and conservative dynamics of a periodic order parameter $\psi\equiv\psi(\mathbf{r})$, respectively. The dynamical equation can be written in terms of the variation of a free energy functional $F$ that, in its simplest form, can be written [30, 35]
$F=\int_{\Omega}{\rm d}\mathbf{r}\bigg{[}\frac{1}{2}\psi\mathcal{L}\psi+\frac{1}{4}\psi^{4}\bigg{]},$ (24)
with $\mathcal{L}=(q_{0}^{2}+\nabla^{2})^{2}-\epsilon$. This differential operator enforces a periodicity in agreement with the first peak in the structure factor (at $q_{0}$) for periodic patterns like stripes or crystalline phases. $\epsilon$ is a phenomenological temperature parameter. The so-called SH equation corresponds to the nonconservative $L^{2}$-gradient flow
@@
Eq. (24) realizes the minimal free energy to model order-disorder phase transition as its minimizers are either constant or periodic fields. As such, its form emerges in different contexts. $F$ can be obtained as the nondimensional form of the free energy describing liquid-solid transition with the order parameter corresponding to the microscopic density field [36]. It may also be derived—upon approximations—from the classical density functional theory [53, 54, 55] as discussed, e.g., in Refs. [56, 57, 37, 58]. In this approach, the differential operator can be obtained as an approximation of an $n$-point correlation function. Other approximations may also be considered, like the multimode SH energy functional [59], effectively reproducing $M$ peaks at wavenumbers $q_{m}$ of the structure factor via a differential operator
@@
$F_{\psi}=\int_{\Omega}{\rm d}\mathbf{r}\bigg{[}A\psi\mathcal{L}_{M}\psi+B\psi^{2}+C\psi^{3}+D\psi^{4}\bigg{]}.$ (28)
@@ 3.2 Amplitude expansion
In the amplitude expansion of the PFC model (APFC) [49], a free energy $F_{\eta}$ dependent on amplitudes can be derived from $F_{\psi}$. For a one-mode approximation of $\psi$ ($M=1$ in $\mathcal{L}_{M}$), this has been obtained via a renormalization group approach or by substituting (29) into the expression of the free energy functional and integrating over the unit cell [62, 63, 64]. Here, with the shortest reciprocal-space vector set to $|\mathbf{k}_{n}|=q_{1}=1$ and focusing on bulk systems where $\psi_{0}(\mathbf{r})$ can be assumed constant ($\psi_{0}(\mathbf{r})=\bar{\psi}$), we consider the amplitude approximation of the multimode SH functional (28) introduced in [65]
@@
with $\mathcal{G}_{n}=\nabla^{2}+2{\rm i}\mathbf{k}_{n}\cdot\nabla$, $g^{\rm s}=\frac{B^{\prime}}{2}\zeta_{2}+\frac{C^{\prime}}{3}\zeta_{3}+\frac{D^{\prime}}{4}\zeta_{4}+E^{\prime}$, $\zeta_{p}$ complex polynomials of order $p$ in the amplitudes depending on the lattice symmetry (see Ref. [49]), and
$\Gamma_{n}=\frac{\prod\limits_{\begin{subarray}{c}m=1\\ q_{m}\neq|\mathbf{k}_{n}|\end{subarray}}^{M}[(q_{m}^{2}-|\mathbf{k}_{n}|^{2})^{2}+b_{m}]}{\prod\limits_{m=2}^{M}[(q_{m}^{2}-q_{1}^{2})^{2}+b_{m}]}$ (31)
@@
4 Elasticity from the SH functional, constants, and characteristic lengths
4.1 Elastic energy with strain-gradient terms
As the displacement field varies over large length scales, elasticity in the SH or PFC model can be described by looking at amplitudes $\{\eta_{j}\}$ [42, 49]. In this context, the elastic energy density, namely the part of the energy depending on the displacement $\mathbf{u}$, is
@@ 4.1 Elastic energy with strain-gradient terms
To proceed further, we consider two widely adopted simplifications, assuming constant amplitude moduli ($|\nabla\varphi_{n}|=0$) and neglecting nonlinear terms. With the former assumption, we obtain
@@
with $\mathbf{U}$ a nonlinear strain tensor and $\mathbf{G}$ encoding gradient-elasticity contributions. Due to the nature of the amplitude functions, this simplifying assumption is exact in relaxed bulk, where the amplitudes are real and constant. However, it is an approximation at defects and interfaces, where the amplitude moduli vary measurably. Next, neglecting higher-order nonlinear terms, we have
$w=A^{\prime}\sum_{n=1}^{N}\Gamma_{n}\varphi_{n}^{2}\left\{4[(\mathbf{k}_{n})_{i}(\mathbf{k}_{n})_{j}(\partial_{i}u_{j})]^{2}+[(\mathbf{k}_{n})_{i}\nabla^{2}u_{i}]^{2}\right\},$ (35)
with the first and second terms in the sum corresponding to linear and strain-gradient elasticity terms, respectively. Due to the presence of higher-order displacement derivatives, it is important to check that this approximated expression of the elastic energy density complies with rotational invariance. Indeed, while the free energy of the system (30) is rotationally invariant by construction [49], and therefore so is the elastic energy density implicitly encoded therein, the approximations made to reach (35) may affect that property. This is, however, not the case, given the form of the double-derivative terms. Explicitly, if the space coordinates undergo a rigid rotation with the associated matrix $\mathbf{R}$, the term $(\mathbf{k}_{n})_{i}\nabla^{2}u_{i}$ transforms as:
$R_{ij}(\mathbf{k}_{n})_{j}\nabla^{2}(R_{ik}u_{k})=\delta_{jk}(\mathbf{k}_{n})_{j}\nabla^{2}u_{k}=(\mathbf{k}_{n})_{j}\nabla^{2}u_{j},$ (36)
i.e., it is rotationally invariant. We refer to Refs. [66, 67] for more detailed discussions. In the following, we look at the expressions of the elastic energy density obtained for specific patterns and lattice symmetries.
4.2 Stripe Phase
Stripe arrangements, e.g., in smectic phases [68, 69, 70, 71], are a class of patterns described by SH and PFC models [30, 35, 72]. The corresponding minimizer of the SH energy functional is well represented by waves with a single reciprocal lattice vector $\mathbf{k}$, namely with $\psi$ described as in Eq. (29) with a single term in the sum. Assuming an arbitrarily oriented wave vector $\mathbf{k}=k_{0}(\cos(\phi),\sin(\phi))$, Eq. (35) reduces to:
${C_{\rm str}}=8A^{\prime}\varphi^{2}k_{0}^{4}\begin{pmatrix}\cos^{4}(\phi)&\cos^{2}(\phi)\sin^{2}(\phi)&\cos^{3}(\phi)\sin(\phi)\\ *&\sin^{4}(\phi)&\cos(\phi)\sin^{3}(\phi)\\ *&*&\cos^{2}(\phi)\sin^{2}(\phi)\end{pmatrix}.$ (38)
Notice that the strain-gradient constants cannot be recast into the characteristic lengths $\ell_{i}$. Stripe phases exhibit significant anisotropy in diffusive processes, such as marked directionality during growth [72, 73, 74]. Here, we obtain that this phase also exhibits elastic anisotropy for linear and strain-gradient elasticity contributions.
4.3 Crystalline phases
For crystalline phases, we find that Eq. (35) may be rewritten as
@@ 4.3 Crystalline phases
given that, for each mode $m$ (i.e., for each family such that $|\mathbf{k}_{n}|=q_{m}$), the symmetry of the corresponding reciprocal lattice vectors (see also below explicit examples) imposes
$\begin{split}\sum_{\begin{subarray}{c}n=1\\ q_{m}=|\mathbf{k}_{n}|\end{subarray}}^{N}(\mathbf{k}_{n})_{i}^{2}=Q_{m},\qquad\text{and}\qquad\sum_{\begin{subarray}{c}n=1\\ q_{m}=|\mathbf{k}_{n}|\end{subarray}}^{N}(\mathbf{k}_{n})_{i}(\mathbf{k}_{n})_{j}=0,\quad\forall i\neq j,\end{split}$ (41)
where $K$ and $Q_{m}$ depend on the specific lattice symmetry and the number of modes considered. To determine GE constants and characteristic lengths, we then need to express the quantity (40) in terms of strain derivatives. In 2D systems, we have that
@@
where we exploited the identity $\partial_{ii}^{2}u_{j}=\partial_{i}(2\varepsilon_{ij}-\partial_{j}u_{i})=2\partial_{i}\varepsilon_{ij}-\partial_{j}\varepsilon_{ii}$. We may then compare terms of $w$ from Eq. (35) to the general form of the elastic energy for isotropic materials (12):
@@
For three-dimensional systems, the expressions of the coefficients $a_{i}$ are analogous. This follows by considering the additional term $(\nabla^{2}u_{z})^{2}$ on the left-hand side of Eq. (42) and proper extension of the right-hand side. The material characteristic length scales, as can be directly deduced from the first and fifth equations in (43), are
@@
with $\ell_{3}=0$. This means that under the assumptions considered in this section (see Eqs. (34) and (35)), the minimal SH energy functional encodes isotropic GE-1. We note that while no centrosymmetric crystal can verify the conditions for isotropic strain-gradient elasticity (due to the Hermann theorem [75]), for cubic materials that are nearly isotropic with respect to the constitutive four-rank tensor $\mathbb{C}$, a Voigt-type averaging procedure may be employed for the sixth-rank constitutive tensor $\mathbb{D}$ to compute effective isotropic GE constants [7].
We consider in the following different crystalline phases: triangular, square, bcc, and fcc. For some representative cases, we analyze different numbers of modes. We report the elastic energy density for the different symmetries/approximations (set by specific choices for $\mathbf{k}_{n}$ with $\min_{n=1}^{N}|\mathbf{k}_{n}|=1$), the corresponding elastic constants, and GE characteristic lengths for small deformations and neglecting nonlinearities (see Table 1 and Fig. 2). The scaling factor $k_{0}$ is introduced such that $a_{0}=2\pi/k_{0}$ is the interatomic distance.
@@
The elastic constants and GE characteristic lengths obtained via Eq. (22) from the elastic energy densities above are reported in Table 1. Note that the constraint $C_{12}=C_{44}$ holds for all lattice symmetries. Interestingly, this corresponds to the Cauchy relation derived in the lattice-theoretical description of constitutive tensors for central force interaction between atoms [76, 7].
@@
4.4 Discussion of the results
From the analysis reported above for crystalline phases, we find that $\ell_{2}/\ell_{1}=\sqrt{3}$ as a result of the general form of the (isotropic) strain-gradient terms; see Eq. (40). Typical values found or adopted in the literature for this ratio are close to one [77, 78]. For a one-mode approximation of $\psi$ (one length of $\mathbf{k}$ considered), length scales are constants. For approximations of $\psi$ considering more than one mode, characteristic lengths depend on $\varphi_{i}$. Figure 1 illustrates the bulk free energy density for the tri-2 (Fig. 1a) and fcc-2 (Fig. 1b) case as a function of $\varphi_{1}$ and $\varphi_{2}$ for selected parameters. The global minima of these free energy densities (blue squares) correspond to the equilibrium phase. Values of $\varphi_{1}$ and $\varphi_{2}$ vary with the free energy parameters as dictated by the energy landscape, so they are not free parameters entering $\alpha$. An illustration of their variation is reported in Fig. 1c (tri-2) and 1d (fcc-2).
Figure 2 reports the values of $\ell_{1,2}$ by varying $B$ in the free energy for a selected parameter range where all the symmetries can be explored for $B<B_{c}$ and $B_{c}$ corresponding to the order-disorder critical point [49]. Here, the effect of changes in $\varphi_{i}$ is then quantified for a specific choice of parameters. We note that a limited change is observed among all the explored ranges ($\lesssim 10\%$). The lattice symmetry affects the values of the GE length scales. At small enough $B$ (low “temperature”), the number of modes considered does not impact the characteristic lengths, whereas a clear deviation is observed at larger $B$ (Fig. 2). In the context of APFC, these approximations realize different models as the choice of modes dictates the variables to solve for in the system (the amplitudes $\eta_{j}$). Instead considering a larger number of modes for the PFC and SH models results in a better estimation of their elastic properties.